A photoelectric surface has a work function of 2.10 eV. Calculate the maximum kinetic energy, in eV, of electrons ejected from this surface by electromagnetic radiation of wavelength 356 nm. (Note: h = 6.63x10^-34 C and c = 3.00x10^8 m/s).
Answer #1
First we want to calculate the amount of energy in the electromagnetic radiation. To do this, we use the formula: E = hv

E = energy of radiation
h = Planck's Constant (6.63x10^-34)
v = frequency of radiation

To find the frequency of the radiation, we use the formula c = fw

c = speed of light (3.0x10^8)
f = frequency of radiation
w = wavelength of radiation (356x10^-9 m)

c = fw
f = c/w = (3.0x10^8)/(356x10^-9)
f = 8.43x10^14 Hz

Now we plug into the equation:

E = (6.63x10^-34)(8.43x10^14) = 5.59x10^-19 J

The conversion factor from joules to eV is 1.602x10^-19 J/eV, so we divide by that number:

5.59x10^-19/1.602x10^-19 = 3.49 eV

We subtract the work function from this to get the amount of kinetic energy:

3.49 eV - 2.10 eV = 1.39 eV of kinetic energy